OpenStudy (anonymous):

find the slope intercept form the line contaning the two points show work . (: (-3,7 ) and ( 6,-8 )

OpenStudy (anonymous):

In order to answer this problem you need to do two basic steps. First you need to find the slope between these two points. The standard form is y=mx+b. If you can find the slope you will get the solution for m. Once you've accomplished this you can plug in the values of one of the given points which should leave you with enough information to solve for b. Here's a basic example. If you have two points (1,2) and (-1,3). We first want to find the slope between these two points. This is given by m = (2-3)/(1-(-1)), which gives us a slope of m = -1/2. In standard form we would have y=(-1/2)x+b. In order to solve for b we simply plugin either given point, say (1,2). So we get 2=(-1/2)(1)+b. We simply solve for b to get 2+(1/2)=b or 5/2=b. We can thus return to the standard form to get the solution for this example: y=(-1/2)x+(5/2).